40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jul 27 15:18:50 2011, kiwamu, Summary, General, schedule 
    Reply  Wed Jul 27 15:42:41 2011, kiwamu, Summary, General, mitues of 40m meeting : Task list 
       Reply  Sun Jul 31 22:48:35 2011, kiwamu, Summary, General, preparation of the vent : status update 
          Reply  Mon Aug 1 20:21:05 2011, kiwamu, Update, Green Locking, Y arm green beam axis : aligned 
             Reply  Tue Aug 2 02:20:09 2011, kiwamu, Update, Green Locking, Y arm green beam axis : done 
Message ID: 5078     Entry time: Sun Jul 31 22:48:35 2011     In reply to: 5048     Reply to this: 5084
Author: kiwamu 
Type: Summary 
Category: General 
Subject: preparation of the vent : status update 

Status update for the vent preparation:

The punchline is : We can not open the chamber on Monday !

 

##### Task List for the vent preparation #####

  (not yet) Low power MC

  (not yet) Measurement of the arm lengths

  (not yet) Alignment of the Y green beam (#5066)

  (not yet) Measurement of the MC spot positions

  (80% done) Estimation of the tolerance of the arm length (#5076)

  (done) Preparation of beam dumps (#5047)

  (done) Health check of shadow sensors and the OSEM damping gain adjustment (#5061)

  (done) Alignment of the incident beam axis (#5073)

  (done) Loss measurement of the arm cavities (#5077)

Quote from #5048

Quote:

The vent will start from 1 st of August ! 

 

ELOG V3.1.3-